# Detecting Misuse of Security APIs: A Systematic Review

### v2

@@ 5. RQ2: Misuses of Security APIs
Through thorough analysis and semantic categorization, we identified 30 distinct misuse types, each assigned a unique identifier (M#). Table 4 lists these misuses and their mappings to relevant security APIs, while Figure 8 illustrates the number of research efforts dedicated to each misuse type. As shown, insecure cryptographic key management (M1), insecure cryptography algorithms (M4), and improper SSL certificate validation (M8) are the most extensively explored misuses in research. Notably, they are also among the most common misuses reported for crypto and SSL APIs. Figure 8 illustrates different findings regarding the most common misuses of security APIs identified in real-world software by reviewed studies. Insecure cryptography algorithms (broken hash) for crypto API, improper certificate validation for SSL/TLS API, lack or misuse of the State parameter (M13) for OAuth API, and lack or misuse of cryptography (M19) for Fingerprint API are the misuses seen with highest frequency in this figure. In the following, we elaborate on these misuse types, discussing their consequences and prevalence in software applications.
@@
M1—Insecure cryptographic key management: M1 is the most extensively studied misuse in the literature, specifically on crypto, SSL/TLS, and Fingerprint APIs (Figure 8). It also ranks among the most common misuses found in various software applications across Android, Windows, and Ubuntu (Figure 8). In the context of crypto and SSL/TLS, M1 often arises from practices like deterministic key generation and poor key derivation, leading to predictable and easily compromised keys. The use of hard-coded, static (constant), expired, previously-used or any kind of predictable keys are among the misuses identified by prior research. Listing LABEL:lst:m1m4 illustrates an example of hard-coded keys identified within an Android app. Keys of inadequate length also increase the possibility of a brute-force attack; for example, RSA keys and ECC keys should be at least 2048 and 224 bits long, respectively (Barker et al., 2009). Additionally, the practice of retaining keys in memory after their intended use creates opportunities for code injection or side-channel attacks, enabling potential key recovery [S13]. Another significant concern arises from insecure key distribution. Using a key agreement protocol that allows a single peer to generate the shared secret without involving the other peers compromises the security of the negotiated key [S13]. Likewise, key exchange without authentication from a trusted entity leads to vulnerabilities to MitM attacks, where malicious servers can impersonate trusted servers and gain access to sensitive information [S15]. Furthermore, ignoring integrity checks is a notable concern. Developers should conduct integrity checks after symmetric key exchange to ensure the integrity of the exchanged keys [S29]. In the context of Fingerprint, M1 often arises from a lack of cryptographic key generation when developers neglect using methods for generating keys to be used for securing fingerprint-based authentication [S38]. Another concern is the use of unlocked keys, where the key remains unlocked, allowing attackers with root access to use it without requiring the user to touch the fingerprint sensor [S38].
@@ 9 SecretKeySpec secretKeySpec = null;
M4—Insecure configurations for encryption: One common misuse of crypto and SSL/TLS APIs is using unsafe modes of operation for encryption, such as Electronic Codebook (ECB). ECB mode encrypts data blocks independently, transforming identical message blocks into identical ciphertext blocks, thus revealing data patterns and compromising confidentiality. To ensure security, it is recommended to use more secure modes, such as Cipher Block Chaining (CBC) or Galois/Counter Mode (GCM). Other instances of unsafe encryption modes include DESede with ECB, DES with CBC3 SHA, AES with CBC and PKCS5Padding, CBC without HMAC, and 3DES with EDE CBC SHA. Additionally, Initialization Vectors (IVs) are crucial in several encryption modes to add entropy to ciphertexts. To ensure the security of cryptographic schemes, IVs must be randomly and properly generated. However, some developers introduce vulnerabilities by using empty, zeroed, hard-coded, static, badly-derived (e.g., deriving from keys or messages), short-length, previously-used, or any kind of predictable IVs. Listing LABEL:lst:m1m4 illustrates an example of hard-coded IVs identified within an Android app. Another parameter that requires secure configuration is the padding scheme, which specifies how to fill the last block of data in encryption. Missing padding or using insecure padding (e.g., PKCS 1-v1.5 for RSA) make it easier for an attacker to launch a padding oracle attack and recover the plaintext.
M5—Insecure configurations for Password-Based Encryption (PBE): Within crypto and SSL/TLS APIs, PBE relies on carefully chosen parameters including salt, password, and iteration count to derive strong encryption keys. Improperly setting these parameters can significantly compromise the security of the derived key. One major misconfiguration of PBE is using an empty, static (constant), short-length (size¡64 bits (Kaliski and Rusch, 2017)), or predictable salt or using a hard-coded, static, weak, NIST-blacklisted, expired, previously-used, or predictable password, which introduces vulnerabilities to brute-force and dictionary attacks. Moreover, developers may prefer to choose small iteration counts (less than 1000 (Kaliski and Rusch, 2017)) to achieve better performance, making it easier for attackers to perform brute-force attacks. Listing LABEL:lst:m5 showcases M5 present in an Android application. The code uses an iteration count based on the password length, falling far below the recommended minimum. Additionally, the salt is generated by hashing the password itself, which introduces another security risk. An attacker gaining access to the salt could potentially recover the password [S1].
@@
M11—Improper error handling: Some developers prioritize functionality over security, leading them to disregard errors, as identified in use cases of the crypto, SSL/TLS, and SafetyNet Attestation APIs. For crypto and SSL/TLS, developers may ignore errors occurring during certificate validation and and simply proceed with normal operations [S24, S34, S43, etc.]. Listing LABEL:lst:m11 exemplifies this by using an empty method for exception handling in an Android app. Likewise, errors may occur during the integrity checks performed by the SafetyNet Attestation API, and disregarding them can result in the failure of the attestation process [S69].
@@ 6 WebView webview;
M17—Insecure OAuth grant types: The security of OAuth transactions highly depends on the choice of grant type. It is essential to avoid using insecure grant types such as implicit for authentication. Implicit grants raise a major security concern because the access token is not bound to the intended RP which enables an attacker to use a user’s access token, issued to the malicious application, to log in as the user on a benign application [S48]. Best current practices recommend using the authorization code flow, which can be protected by PKCE [S54].
@@
M30—Performing attestation only at first launch: SafetyNet Attestation should be consistently performed during an application life cycle, specifically when launching or handling sensitive information. However, some developers only perform SafetyNet Attestation during the first launch, leaving the application vulnerable to tampering [S69]. Attackers could exploit this by launching the application once in a non-tampered state, then tamper the device or application later without being detected, as SafetyNet Attestation will not be performed anymore.
@@ 6.1. Types of Modeling Input
Figure 10. Distribution of misuse detection techniques adopted by primary studies over (a) modeling input types (b) testing input types (c) output types (d) automation mode
@@ 6.2. Types of Testing Input
Various input types are utilized to analyze an application for the presence of misuse, including source code, binary code, intermediate representation, or applications, depending on the analysis algorithm employed (Section 6.5 details analysis algorithms). The most common approach is analyzing source code employed in 33 studies. Some misuse detection tools are developed as IDE plugins to provide real-time feedback as developers write code. For example, FixDroid [S5] was introduced as a plugin for Android Studio, or CogniCrypt [S3] as an Eclipse plugin. However, source code analysis has limitations. Existing approaches are often limited to specific programming languages. A study found developers frequently use C for cryptography in MicroPython projects, highlighting the need for multi-language tools [S20]. Additionally, source code analysis is typically restricted to open-source applications. Some studies [S14, S23, S26] decompiled binary code to gain access to source codes when the original source code is unavailable. However, obfuscation techniques, that modify class files to protect the source code, present challenges to successful decompilation.
In the absence of source code, 25 studies employed binary code analysis. Binary files are typically packaged with other resources such as images and configuration files into application files. Thus, misuse detection techniques typically require reverse engineering to decode an APK file and disassemble it into binary files, ready for analysis. Disassembling is typically faster than decompiling and is not affected by obfuscation. Common program analysis frameworks used by primary studies to extract binary files from application files include SOOT (Lam et al., 2011), APKTool (Connor and Ryszard, 2010), and Androguard (Desnos, 2012).
@@
Meanwhile, 22 studies [e.g., S6, S8, S10, S13, S14] detected misuse by executing applications and analyzing their runtime, eliminating the need for accessing source or binary code. However, this method is resource-intensive and time-consuming due to the deployment and execution phases. Figure 10.b illustrates the distribution of testing input types across the reviewed studies. Notably, several studies employed multiple input types for analysis, which were categorized under all relevant types. For instance, some studies [e.g., S19, S24] examined potential misuses by first analyzing binary code and then detected exploitable misuses at runtime by executing the applications.
@@ 6.3. Output Types
Common ways to support developers in addressing misuses are issuing misuse alarms (69 studies), performing source attribution (4 studies), generating security advice (4 studies), and assisting with fixing (4 studies) (Figure 10.c). In the following, we elaborate on output types other than misuse alarms which were exclusively explored in the context of crypto and SSL/TLs APIs by the literature.
Source attribution: The objective of source attribution is to attribute whether a misuse originates from an application code or a third-party library. It is beneficial for developers to identify libraries with misuses and avoid using them, and for researchers to avoid over-counting misuses by identifying those that stem from libraries [S12]. Three studies investigated the primary source of crypto and SSL/TLS misuse and, interestingly, showed that third-party libraries are the major reason for misuse in Android applications [S1, S12] and popular Python projects in GitHub [S20]. Another study [S35] evaluated a large database of third-party libraries and found that crypto misuses are very common among widely used advertising libraries. It further identified affected Android apps through third-party library detection.
@@
Repair: Studies exploring security advice and fix suggestions often lack the ability to provide customized fixes for a given vulnerable program. Only 4 studies [S2, S16, S17, S46] proposed methods for generating such fixes for crypto and SSL/TLS misuses. One simple approach involves manually crafting patch templates, as performed by CDRep [S16] and CryptoTutor [S17]. However, these templates have limited coverage for diverse misuse variations. In contrast, Seader [S2] and VuRLE [S46] automatically generated customized fixes from code examples. They compared ASTs of a given vulnerable code and its corresponding repaired code to extract the edit operations necessary to transform a vulnerable program into a secure one. Using edit operations, Seader generated an abstract fix for a vulnerable program including abstract variables which were replaced by concrete variables to customize fixes. On the other hand, VuRLE customized the edit patterns for a specific vulnerable program and applied the customized changes to repair vulnerabilities.
@@ 6.5.1. Heuristic-based
A. Pattern Types: API misuses can be detected using two types of patterns: normal (12 studies) or misuse (54 studies). Misuse detection through normal patterns involves modeling correct API use and identifying any deviations from them as misuse. While APIs can be misused in numerous ways, only a small subset corresponds to proper usage. Thus, this approach can identify a wide range of misuse types through limited normal patterns. However, this approach suffers from producing high false alarms when it fails to model normal patterns thoroughly. An alternative solution is to model incorrect API uses and identify any matches with them as misuse. However, predicting all possible ways that a developer could misuse an API is a challenging task, so this approach may not capture all misuses. Nevertheless, most of the studies in our review rely on misuse patterns to avoid high false alarms in approaches based on normal patterns.
@@
(i) Language-based templates rely on a syntax-based representation of patterns. CrySL [S11], for instance, is a language designed for crypto experts to specify the normal use of crypto APIs. Several studies [S3, S11, S36, S40, S58, S61] used CrySL to detect crypto misuses. Meta-CrySL [S55] is an extension of CrySL that helps manage variations in the API and security standards specified within CrySL. Another study [S56] introduced a formal model for security annotations that describe properties ensuring the secure usage of WebCrypto API. Furthermore, an anti-protocol language was introduced to describe common misuse patterns for OAuth API [S30].
(ii) Graph-based templates represent key elements used while interacting with APIs as nodes and the correlations between these elements as edges. SSLint [S9], for instance, models the proper use of SSL API based on the program dependency graph representing critical API call sites, variables, parameters, and conditions.
@@
(iv) Code-based templates: Crafting language-, graph-, and FSM-based templates requires substantial domain knowledge and manual effort to specify the critical elements for API usage, their correlation, and modeling them as templates. An alternative solution is to automatically infer templates from sample security API use cases within source code. In our review, one study [S2] used (insecure, secure) code pairs from prior research and compared their ASTs to identify edit operations. Vulnerability and repair templates were then generated based on a data-dependency analysis of ASTs and variable abstraction. The vulnerability template was used to detect misuse through pattern matching, while the repair template was used to generate customized fixes. Another study [S46] analyzed commits from 48 GitHub applications to manually identify misuses. ASTs were then used to identify required edit operations for fixes, which were clustered later based on their similarities. Finally, each cluster was generalized to vulnerability and repair templates to be used for detecting and fixing misuses. While code templates enable automatic pattern generation, they are inherently limited to known misuses reflected in existing code.
Table 5 summarizes descriptions, strengths and weaknesses of pattern types and representation models and Figure 12.a shows their distribution in primary studies.
@@
Normal Patterns are inferred from correct uses of APIs and any violation of these patterns is considered as misuse. $\bullet$ Limited number of patterns
@@
Misuse Patterns are inferred from incorrect uses of APIs and any matches with these patterns are considered as misuses. $\bullet$ Incomplete specification does not result in false alarms $\bullet$ Difficult to capture all possible patterns
@@
Template Patterns are abstracted via a higher level template. $\bullet$ Easier to extend to new misuses $\bullet$ Difficult to design a template from instances
@@
C. Program Analysis Techniques: Our review identified three categories for dividing program analysis techniques based on their reliance on code execution: (i) static analysis, (ii) dynamic analysis and (iii) hybrid analysis. In the following, we investigate these methods and their adoption in the existing literature.
(i) Static Analysis: Static analysis involves examining (recovered) source code, binary code, or an intermediate representation of binary code without executing the application. This approach is also known as white-box testing as it requires application code or its implementation details to identify misuses. Static analysis offers significant advantages: it is resource- and time-efficient, and it can achieve high code coverage. Additionally, developers can integrate static analysis tools into their daily workflow to identify misuses early in the development process. However, static analysis also has limitations, such as high false alarm rates caused by infeasible misuses (that never occur at runtime) and failure to capture runtime misuses.
@@
To achieve more efficient misuse detection, several studies applied program slicing, which simplifies the complexity of a program by removing parts of the code that are irrelevant to a specific analysis or task (Xu et al., 2005). This is accomplished by computing a set of program statements that affect (backward slicing) or are affected by (forward slicing) a given slicing criterion, which is typically an API parameter, based on data flow [S1]. For instance, CryptoTutor [S17] applied inter-procedural data flow analysis and program slicing to detect crypto misuses in Java code. Similarly, CryptoLint [S4] used inter-procedural backward slicing to track flows between crypto parameters and operations. Later, BinSight [S12] and CDRep [S16] leveraged CryptoLint to examine the current state of crypto API usage in Android applications, with additional efforts towards source attribution [S12] and repair [S16]. Amandroid [S44] also applied inter-procedural data flow analysis to assess the security state of Android apps in terms of data leaks, data injection, and improper use of crypto and SSL/TLS APIs. Program slicing was also used in the context of Fingerprint API to classify applications into different security levels based on the state of their API use. This involved performing inter-procedural backward slicing to extract API parameters as features in a rule-based classification system [S37].
@@
The passive dynamic analysis examines execution logs, the memory state of a program, or network traces to gain insight into its behavior. Considering the observed data, we have categorized passive dynamic approaches into log, memory, and network analysis. Log analysis involves collecting runtime information and execution traces and performing offline analysis after the execution is completed. While the offline analysis does not affect the application’s performance [S10], it can generate large log files, creating an I/O bottleneck slowdown [S13]. For example, one study [S10] examined logs that record parameters relevant to crypto API calls to find matches with some misuse patterns. In our review, log analysis was performed by a few studies to detect misuse of security APIs. Memory analysis was also utilized by some studies for misuse detection. For example, K-Hunt [S13] tracks memory buffers that store encryption keys to verify whether keys are generated and transmitted securely. It started with a lightweight dynamic analysis to gather runtime information required to locate memory buffers where crypto keys were stored. Meanwhile, several studies adopted network analysis approach to detect misuses of APIs such as SSL/TLS and OAuth. One study [S52] evaluated the implementation of CSRF protection in OAuth transactions by checking the presence or absence of a state variable in URLs.
In addition, some studies performed active dynamic analysis to verify the results obtained from network analysis. Active dynamic analysis can include a range of techniques, such as penetration testing which simulates a real-world attack on a running application to identify any misuses that could be exploited. It is considered as the most effective approach to uncover exploitable misuses and avoid false alarms. One study [S50] manually analyzed the HTTP messages to capture the information flow of SSO credentials and detect potential misuse of OAuth. It further designed exploits to prevent manual inspection errors. Similarly, another study [S49] performed network analysis followed by examining the feasibility of a CSRF attack to uncover exploitable misuses of OAuth.

### v3

@@ 5. RQ2: Misuses of Security APIs
Through thorough analysis and semantic categorization, we identified 30 distinct misuse types, each assigned a unique identifier (M#). Table 4 lists these misuses and their mappings to relevant security APIs, while Figure 8 illustrates the number of research efforts dedicated to each misuse type. As shown, insecure cryptographic key management (M1), insecure cryptography algorithms (M4), and improper SSL certificate validation (M8) are the most extensively explored misuses in research. Notably, they are also among the most common misuses reported for crypto and SSL APIs. Figure 8 illustrates different findings regarding the most common misuses of security APIs identified in real-world software by reviewed studies. Insecure cryptography algorithms (broken hash) for crypto API, improper certificate validation for SSL/TLS API, lack or misuse of the State parameter (M13) for OAuth API, and lack or misuse of cryptography (M19) for Fingerprint API are the misuses seen with the highest frequency in this figure. In the following, we elaborate on these misuse types, discussing their consequences and prevalence in software applications.
@@
M1—Insecure cryptographic key management: M1 is the most extensively studied misuse in the literature, specifically on crypto, SSL/TLS, and Fingerprint APIs (Figure 8). It also ranks among the most common misuses found in various software applications across Android, Windows, and Ubuntu (Figure 8). In the context of crypto and SSL/TLS, M1 often arises from practices like deterministic key generation and poor key derivation, leading to predictable and easily compromised keys. The use of hard-coded, static (constant), expired, previously-used or any kind of predictable keys is among the misuses identified by prior research. Listing LABEL:lst:m1m4 illustrates an example of hard-coded keys identified within an Android app. Keys of inadequate length also increase the possibility of a brute-force attack; for example, RSA keys and ECC keys should be at least 2048 and 224 bits long, respectively (Barker et al., 2009). Additionally, the practice of retaining keys in memory after their intended use creates opportunities for code injection or side-channel attacks, enabling potential key recovery [S13]. Another significant concern arises from insecure key distribution. Using a key agreement protocol that allows a single peer to generate the shared secret without involving the other peers compromises the security of the negotiated key [S13]. Likewise, key exchange without authentication from a trusted entity leads to vulnerabilities to MitM attacks, where malicious servers can impersonate trusted servers and gain access to sensitive information [S15]. Furthermore, ignoring integrity checks is a notable concern. Developers should conduct integrity checks after symmetric key exchange to ensure the integrity of the exchanged keys [S29]. In the context of Fingerprint, M1 often arises from a lack of cryptographic key generation when developers neglect using methods for generating keys to be used for securing fingerprint-based authentication [S38]. Another concern is the use of unlocked keys, where the key remains unlocked, allowing attackers with root access to use it without requiring the user to touch the fingerprint sensor [S38].
@@ 9 SecretKeySpec secretKeySpec = null;
M4—Insecure configurations for encryption: One common misuse of crypto and SSL/TLS APIs is using unsafe modes of operation for encryption, such as Electronic Codebook (ECB). ECB mode encrypts data blocks independently, transforming identical message blocks into identical ciphertext blocks, thus revealing data patterns and compromising confidentiality. To ensure security, it is recommended to use more secure modes, such as Cipher Block Chaining (CBC) or Galois/Counter Mode (GCM). Other instances of unsafe encryption modes include DESede with ECB, DES with CBC3 SHA, AES with CBC and PKCS5Padding, CBC without HMAC, and 3DES with EDE CBC SHA. Additionally, Initialization Vectors (IVs) are crucial in several encryption modes to add entropy to ciphertexts. To ensure the security of cryptographic schemes, IVs must be randomly and properly generated. However, some developers introduce vulnerabilities by using empty, zeroed, hard-coded, static, badly-derived (e.g., deriving from keys or messages), short-length, previously-used, or any kind of predictable IVs. Listing LABEL:lst:m1m4 illustrates an example of hard-coded IVs identified within an Android app. Another parameter that requires secure configuration is the padding scheme, which specifies how to fill the last block of data in encryption. Missing padding or using insecure padding (e.g., PKCS 1-v1.5 for RSA) makes it easier for an attacker to launch a padding oracle attack and recover the plaintext.
M5—Insecure configurations for Password-Based Encryption (PBE): Within crypto and SSL/TLS APIs, PBE relies on carefully chosen parameters including salt, password, and iteration count to derive strong encryption keys. Improperly setting these parameters can significantly compromise the security of the derived key. Common misconfigurations of PBE include using an empty, static (constant), short-length (size¡64 bits (Kaliski and Rusch, 2017)), or predictable salt or using a hard-coded, static, weak, NIST-blacklisted, expired, previously-used, or predictable password, which introduces vulnerabilities to brute-force and dictionary attacks. Moreover, developers may prefer to choose small iteration counts (less than 1000 (Kaliski and Rusch, 2017)) to achieve better performance, making it easier for attackers to perform brute-force attacks. Listing LABEL:lst:m5 showcases M5 present in an Android application. The code uses an iteration count based on the password length, falling far below the recommended minimum. Additionally, the salt is generated by hashing the password itself, which introduces another security risk. An attacker gaining access to the salt could potentially recover the password [S1].
@@
M11—Improper error handling: Some developers prioritize functionality over security, leading them to disregard errors, as identified in use cases of the crypto, SSL/TLS, and SafetyNet Attestation APIs. For crypto and SSL/TLS, developers may ignore errors occurring during certificate validation and simply proceed with normal operations [S24, S34, S43, etc.]. Listing LABEL:lst:m11 exemplifies this by using an empty method for exception handling in an Android app. Likewise, errors may occur during the integrity checks performed by the SafetyNet Attestation API, and disregarding them can result in the failure of the attestation process [S69].
@@ 6 WebView webview;
M17—Insecure OAuth grant types: The security of OAuth transactions highly depends on the choice of grant type. It is essential to avoid using insecure grant types, such as implicit for authentication. Implicit grants raise a major security concern because the access token is not bound to the intended RP, which enables an attacker to use a user’s access token, issued to the malicious application, to log in as the user on a benign application [S48]. Best current practices recommend using the authorization code flow, which can be protected by PKCE [S54].
@@
M30—Performing attestation only at first launch: SafetyNet Attestation should be consistently performed during an application life cycle, specifically when launching or handling sensitive information. However, some developers only perform SafetyNet Attestation during the first launch, leaving the application vulnerable to tampering [S69]. An attacker could exploit this by initially launching the application in a non-tampered state, then subsequently tampering with the device or application without detection, as SafetyNet Attestation would no longer be performed.
@@ 6.1. Types of Modeling Input
Figure 10. Distribution of misuse detection techniques adopted by primary studies over (a) modeling input types, (b) testing input types, (c) output types, (d) automation mode
@@ 6.2. Types of Testing Input
Various input types are utilized to analyze an application for the presence of misuse, including source code, binary code, intermediate representation, or applications, depending on the analysis algorithm employed (Section 6.5 details analysis algorithms). The most common approach is analyzing the source code, which was employed in 33 studies. Some misuse detection tools are developed as IDE plugins to provide real-time feedback as developers write code. For example, FixDroid [S5] was introduced as a plugin for Android Studio, or CogniCrypt [S3] as an Eclipse plugin. However, source code analysis has limitations. Existing approaches are often limited to specific programming languages. A study found developers frequently use C for cryptography in MicroPython projects, highlighting the need for multi-language tools [S20]. Additionally, source code analysis is typically restricted to open-source applications. Some studies [S14, S23, S26] decompiled binary code to gain access to source code when the source code is unavailable. However, obfuscation techniques that modify class files to protect the source code present challenges to successful decompilation.
In the absence of source code, 25 studies employed binary code analysis. Binary files are typically packaged with other resources, such as images and configuration files, into application files. Thus, misuse detection techniques typically require reverse engineering to decode an APK file and disassemble it into binary files, ready for analysis. Disassembling is typically faster than decompiling and is not affected by obfuscation. Common program analysis frameworks used by primary studies to extract binary files from application files include SOOT (Lam et al., 2011), APKTool (Connor and Ryszard, 2010), and Androguard (Desnos, 2012).
@@
Meanwhile, 22 studies [e.g., S6, S8, S10, S13, S14] detected misuse by executing applications and analyzing their runtime, eliminating the need for accessing source or binary code. However, this method is resource-intensive and time-consuming due to the deployment and execution phases. Figure 10.b illustrates the distribution of testing input types across the reviewed studies. Notably, several studies employed multiple input types for analysis, which were categorized under all relevant types. For instance, some studies [e.g., S19, S24] examined potential misuses by analyzing binary code and then detected exploitable misuses at runtime by executing the applications.
@@ 6.3. Output Types
Common ways to support developers in addressing misuses are issuing misuse alarms (69 studies), performing source attribution (4 studies), generating security advice (4 studies), and assisting with fixing (4 studies) (Figure 10.c). In the following, we elaborate on output types other than misuse alarms, which were exclusively explored in the context of crypto and SSL/TLS APIs by the literature.
Source attribution: The objective of source attribution is to determine whether a misuse originates from an application code or a third-party library. It is beneficial for developers to identify libraries with misuses and avoid using them, and for researchers to avoid over-counting misuses by identifying those that stem from libraries [S12]. Three studies investigated the primary source of crypto and SSL/TLS misuse and, interestingly, showed that third-party libraries are the major reason for misuse in Android applications [S1, S12] and popular Python projects in GitHub [S20]. Another study [S35] evaluated a large database of third-party libraries and found that crypto misuses are very common among widely used advertising libraries. It further identified affected Android apps through third-party library detection.
@@
Repair: Studies exploring security advice and fix suggestions often lack the ability to provide customized fixes for a given vulnerable program. Only 4 studies [S2, S16, S17, S46] proposed methods for generating such fixes for crypto and SSL/TLS misuses. One simple approach involves manually crafting patch templates, as performed by CDRep [S16] and CryptoTutor [S17]. However, these templates have limited coverage for diverse misuse variations. In contrast, Seader [S2] and VuRLE [S46] automatically generated customized fixes from code examples. They compared ASTs of a given vulnerable code and its corresponding repaired code to extract the edit operations necessary to transform a vulnerable program into a secure one. Using edit operations, Seader generated an abstract fix for a vulnerable program, including abstract variables that were replaced by concrete variables to customize fixes. On the other hand, VuRLE customized the edit patterns for a specific vulnerable program and applied the customized changes to repair vulnerabilities.
@@ 6.5.1. Heuristic-based
A. Pattern Type: API misuses can be detected using two types of patterns: normal (12 studies) or misuse (54 studies). Misuse detection through normal patterns involves modeling correct API use and identifying any deviations from it as misuse. While APIs can be misused in numerous ways, only a small subset corresponds to proper usage. Thus, this approach can identify a wide range of misuse types through limited normal patterns. However, this approach suffers from producing high false alarms when it fails to model normal patterns thoroughly. An alternative solution is to model incorrect API uses and identify any matches with them as misuse. However, predicting all possible ways that a developer could misuse an API is a challenging task, so this approach may not capture all misuses. Nevertheless, most of the studies in our review rely on misuse patterns to avoid high false alarms in approaches based on normal patterns.
@@
(i) Language-based templates rely on a syntax-based representation of patterns. CrySL [S11], for instance, is a language designed for crypto experts to specify the normal use of crypto APIs. Several studies [S3, S11, S36, S40, S58, S61] used CrySL to detect crypto misuses. Meta-CrySL [S55] is an extension of CrySL that helps manage variations in the API and security standards specified within CrySL. Another study [S56] introduced a formal model for security annotations that describe properties ensuring the secure usage of the WebCrypto API. Furthermore, an anti-protocol language was introduced to describe common misuse patterns for the OAuth API [S30].
(ii) Graph-based templates represent key elements used while interacting with APIs as nodes and the correlations between these elements as edges. SSLint [S9], for instance, models the proper use of the SSL/TLS API based on the program dependency graph representing critical API call sites, variables, parameters, and conditions.
@@
(iv) Code-based templates: Crafting language-, graph-, and FSM-based templates requires substantial domain knowledge and manual effort to specify the critical elements for API usage, their correlation, and modeling them as templates. An alternative solution is to automatically infer templates from sample security API use cases within the source code. In our review, one study [S2] used (insecure, secure) code pairs from prior research and compared their ASTs to identify edit operations. Vulnerability and repair templates were then generated based on a data-dependency analysis of ASTs and variable abstraction. The vulnerability template was used to detect misuse through pattern matching, while the repair template was used to generate customized fixes. Another study [S46] analyzed commits from 48 GitHub applications to manually identify misuses. ASTs were then used to identify required edit operations for fixes, which were clustered later based on their similarities. Finally, each cluster was generalized to vulnerability and repair templates to be used for detecting and fixing misuses. While code templates enable automatic pattern generation, they are inherently limited to known misuses reflected in existing code.
Table 5 summarizes descriptions, strengths, and weaknesses of pattern types and representation models, and Figure 12.a shows their distribution in primary studies.
@@
Normal Patterns are inferred from correct uses of APIs, and any violation of these patterns is considered as misuse. $\bullet$ Limited number of patterns
@@
Misuse Patterns are inferred from incorrect uses of APIs, and any matches with these patterns are considered as misuses. $\bullet$ Incomplete specification does not result in false alarms $\bullet$ Difficult to capture all possible patterns
@@
Template Patterns are abstracted via a higher-level template. $\bullet$ Easier to extend to new misuses $\bullet$ Difficult to design a template from instances
@@
C. Program Analysis Techniques: Our review identified three categories for dividing program analysis techniques based on their reliance on code execution: (i) static analysis, (ii) dynamic analysis, and (iii) hybrid analysis. In the following, we investigate these methods and their adoption in the existing literature.
(i) Static Analysis: Static analysis involves examining (recovered) source code, binary code, or an intermediate representation of binary code without executing the application. This approach is also known as white-box testing as it requires application code or its implementation details to identify misuses. Static analysis offers significant advantages: it is resource- and time-efficient, and it can achieve high code coverage. Additionally, developers can integrate static analysis tools into their daily workflow to identify misuse early in the development process. However, static analysis also has limitations, such as high false alarm rates caused by infeasible misuses (that never occur at runtime) and failure to capture runtime misuses.
@@
To achieve more efficient misuse detection, several studies applied program slicing, which simplifies the complexity of a program by removing parts of the code that are irrelevant to a specific analysis or task (Xu et al., 2005). This is accomplished by computing a set of program statements that affect (backward slicing) or are affected by (forward slicing) a given slicing criterion, which is typically an API parameter, based on data flow [S1]. For instance, CryptoTutor [S17] applied inter-procedural data flow analysis and program slicing to detect crypto misuses in Java code. Similarly, CryptoLint [S4] used inter-procedural backward slicing to track flows between crypto parameters and operations. Later, BinSight [S12] and CDRep [S16] leveraged CryptoLint to examine the current state of crypto API usage in Android applications, with additional efforts towards source attribution [S12] and repair [S16]. Amandroid [S44] also applied inter-procedural data flow analysis to assess the security state of Android apps in terms of data leaks, data injection, and improper use of crypto and SSL/TLS APIs. Program slicing was also used in the context of the Fingerprint API to classify applications into different security levels based on the state of their API use. This involved performing inter-procedural backward slicing to extract API parameters as features in a rule-based classification system [S37].
@@
The passive dynamic analysis examines execution logs, the memory state of a program, or network traces to gain insight into its behavior. Considering the observed data, we have categorized passive dynamic approaches into log, memory, and network analysis. Log analysis involves collecting runtime information and execution traces and performing offline analysis after the execution is completed. While the offline analysis does not affect the application’s performance [S10], it can generate large log files, creating an I/O bottleneck slowdown [S13]. In our review, a few studies performed log analysis to detect misuse of security APIs. For example, one study [S10] examined logs that record parameters relevant to crypto API calls to find matches with some misuse patterns. Memory analysis was also utilized by some studies for misuse detection. For example, K-Hunt [S13] tracks memory buffers that store encryption keys to verify whether keys are generated and transmitted securely. It started with a lightweight dynamic analysis to gather runtime information required to locate memory buffers where crypto keys were stored. Meanwhile, several studies adopted network analysis to detect misuses of APIs such as SSL/TLS and OAuth. One study [S52] evaluated the implementation of CSRF protection in OAuth transactions by checking the presence or absence of a state variable in URLs.
In addition, some studies performed active dynamic analysis to verify the results obtained from network analysis. Active dynamic analysis can include a range of techniques, such as penetration testing, which simulates a real-world attack on a running application to identify any misuses that could be exploited. It is considered the most effective approach to uncover exploitable misuses and avoid false alarms. One study [S50] manually analyzed the HTTP messages to capture the information flow of SSO credentials and detect potential misuse of OAuth. It further designed exploits to prevent manual inspection errors. Similarly, another study [S49] performed network analysis, followed by examining the feasibility of a CSRF attack to uncover exploitable misuses of OAuth.